Please help me I'm so confused, I'm taking a chem class and I'm just not getting conversions. Heres one of the problems I don't get. A room has dimensions of 10.0ftx20.0ftx8.00ft. What is the volume of the room in yd to the exponent 3 on yards.

Answer by Alan3354(69443)

A room has dimensions of 10.0ftx20.0ftx8.00ft. What is the volume of the room in yd to the exponent 3 on yards.
------------
1 yd = 3 feet (another clever English unit)
10 by 20 by 8 feet = (10/3)*(20/3)*(8/3) yds
= 10*20*8/27
= 1600/27
=~ 59.259 cubic yds